How do i solve 49 times 19... Mentally?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

How do i solve 49 times 19... Mentally?
-------
49*19 = (50-1)(20-1) = 1000-50-20+1 = 931
==================
Cheers,
Stan H.

Answer by Alan3354(69443)

How do i solve 49 times 19... Mentally?
----------------
There are many ways for this one.
---
= 19*7 then *7
= 19*50 then -19 the best for this one, imo
---------------
Another method, not so attractive for this one is
Get the average = (49 + 19)/2 = 34
Diff between ave and numbers = 15
49*19 = avg^2 - diff^2 = 1156 - 225
This works well for a lot of them, not so good for this one.
eg, 37*39 = 1443 easily (if you know what 38^2 is)